OpenStudy (anonymous):

Perform the division.-36r^4+40r^324r^2+23r+7/-4r

OpenStudy (campbell_st):

I'll assume that the term in r^2 is being added. well it can be simplifed to \[\frac{-36r^4}{-4r} + \frac{40r^3}{-4r} + \frac{24r^2}{-4r} + \frac{23r}{-4r} + \frac{7}{-4r}\] this will allow the expression to be simplified term by term. The last term will be the remainder. and the constant term will be a fraction. hope this helps.

OpenStudy (anonymous):

that helped a lot thank you!